• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hyde

    Log entry time 23:48:11 on March14,2009

    Entry number 266159

    keyword=Evening Shift Summary

    Shift Summary
    Evening Shift Saturday 14 March 2009

    Exp E06-014
    Will start Kin 4H

    16:00
    Target Heater replacement

    17:30
    Larry Cardman in Hall for tour

    17;45
    Beam restore
    Put Half Wave Plate in
    BeO target
    SPOT++, BPM OK, Harp scan,
    N2 Raster study 2x2 to 6x6

    18:06 Production Data on N2 Reference cell
    Runs
    HRS-L # 20685
    BB # 2154

    18:20 BB WC HV trip, beam trip
    18:35 continue N2
    Runs
    HRS-L 20686
    BB 2155
    18;55 End runs
    Switch to 3He NMR measurement
    Target temperature alarms
    Target Polarization 48%

    19:26 3He target 15 muA
    Runs
    HRS-L 20687
    BB 2156
    20:19
    End run BB 2156 6.5M, let HRS-L 20687 continue
    20:20 Start BB run 2157
    20:23 Brad: end BB run, reboot ROC5 TDC problem

    20:28 Start BB run 2158 HRS-L 20687 continues
    21:18 End runs
    HRS-L 20687 0.099 C
    BB 2158

    21:19 NMR: Target Pol = 54%
    21:20 Start Runs
    HRS-L 20688
    BB 2159
    21:35 BB DC HV L14.7(?) trip, beam trip
    cycle HV
    end BB run 2159
    21:38 Start BB run 2160
    MCC Crew chief says they observed Orbit Shift, but not sure of cause

    23:20 End Runs
    HRS-L 20688 0.0998 C
    BB 2161
    NMR 56.6 %

    23:22 Start Runs
    HRS-L 20689
    BB 2162

    0.2 C in 4hr